HYDROGEN BONDING EFFECTS ON THE INFRARED INTENSITIES OF HALIDE-WATER COM-PLEXES

\underline{SAMANTHA HORVATH} and ANNE B. McCOY, Department of Chemistry, The Ohio State University, Columbus, OH 43210; JOSEPH R. ROSCIOLI and MARK A. JOHNSON, Sterling Chemistry Laboratories, Yale University, New Haven, CT 06520.

The implications of H-bonding in halide-water complexes and its effects on infrared spectra have been of long standing interest, with studies focusing mainly on the OH stretching region. Until recently only the low-energy portion of the X^{−}(H_{2}O) spectra had been probed experimentally (X = F, Cl, Br).^{a}^{,}^{b} An interesting feature in this low-energy region, especially noticeable in the Cl^{−}(H_{2}O) spectrum,^{b} is the large intensity of the overtone for the out-of-plane bending mode. The potential and dipole moment surfaces along this internal coordinate are calculated at the MP2 level of theory (aug-cc-pVTZ basis), and the X^{−} vibrational wave functions and energies are the result of one-dimensional variational calculations.^{c} Our results reproduce the experimental findings very well, and we determine that the overtone intensities reflect the large shifts in charge distribution along the heavy-atom axis when the hydrogen moves out of the plane. This method, although only one-dimensional, shows excellent agreement with experiment and accurately explains the essential physics that give rise to these spectral features.

aG. M. Chaban, S. S. Xantheas, and R. B. Gerber, J. Phys. Chem. A \underline{107}, 4952 (2003).

bE. G. Diken et al., J. Phys. Chem. A \underline{109}, 571 (2005).

cD. T. Colbert and W. H. Miller, J. Chem. Phys. \underline{96}, 1982 (1992).